40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 15 10:52:38 2019, gautam, Update, SUS, EY door opened 
    Reply  Tue Jan 15 15:49:47 2019, gautam, Update, SUS, EY door opened EY_OSEMs.png
       Reply  Sat Jan 19 15:33:18 2019, gautam, Update, SUS, ETMY OSEMs faulty Screen_Shot_2019-01-19_at_3.32.35_PM.pngETMY_sensors_1231832635.pdf
          Reply  Tue Jan 22 20:36:53 2019, gautam, Update, SUS, ETMY OSEMs faulty 
             Reply  Wed Jan 23 12:39:18 2019, gautam, Update, SUS, EY chamber work 6x
                Reply  Wed Jan 23 23:12:44 2019, gautam, Update, SUS, Prep for FC cleaning IY_level_before.pdf
                   Reply  Thu Jan 24 15:32:31 2019, gautam, Update, SUS, Y arm cavity side first contact applied Yarm_FC.pdfOSEMfilter.png
                      Reply  Tue Jan 29 22:12:40 2019, gautam, Update, SUS, Alignment prep 
                         Reply  Wed Jan 30 11:54:24 2019, gautam, Update, SUS, More alignment prep IMG_5958.JPGIMG_5962.JPG
                            Reply  Wed Jan 30 19:25:40 2019, gautam, Update, SUS, Xarm cavity alignment Xarm.pngXarm_IR.png
                               Reply  Fri Feb 1 01:24:06 2019, gautam, Update, SUS, Almost ready for pumpdown tomorrow 
                                  Reply  Fri Feb 1 13:16:50 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                     Reply  Fri Feb 1 21:52:57 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                  Reply  Fri Feb 1 14:44:14 2019, gautam, Update, SUS, Y arm FC cleaning and reinstall IMG_5974.JPGIMG_5986.JPGIMG_5992.JPG
    Reply  Wed Jan 16 16:25:25 2019, gautam, Update, SUS, Yarm locked Yarm_locked.png
       Reply  Fri Jan 18 21:34:18 2019, gautam, Update, SUS, Unused optic on EY table IMG_5408.JPG
          Reply  Sat Jan 19 05:07:45 2019, Koji, Update, SUS, Unused optic on EY table 
Message ID: 14409     Entry time: Sat Jan 19 15:33:18 2019     In reply to: 14401     Reply to this: 14411
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY OSEMs faulty 

After diagnosis with the tester box, as I suspected, the fully open DC voltages on the two problematic channels, LL and UR, were restored once I replaced the LM6321 ICs in those two channel paths. However, I've been puzzled by the inability to turn on the Oplev loops on ETMY. Furthermore, the DC bias voltages required to get ETMY to line up with the cavity axis seemed excessively large, particularly since we seemed to have improved the table levelling.

I suspected that the problem with the OSEMs hasn't been fully resolved, so on Thursday night, I turned off the ETMY watchdog, kicked the optic, and let it ringdown. Then I looked at the time-series (Attachment #1) and spectra (Attachment #2) of the ringdowns. Clearly, the LL channel seems to saturate at the lower end at ~440 counts. Moreover, in the time domain, it looks like the other channels see the ringdown cleanly, but I don't see the various suspension eigenmodes in any of the sensor signals. I confirmed that all the magnets are still attached to the optic, and that the EQ stops are well clear of the optic, so I'm inclined to think that this behavior is due to an electrical fault rather than a mechanical one.

For now, I'll start by repeating the ringdown with a switched out Satellite Box (SRM) and see if that fixes the problem. 

Quote:

While restoring OSEMs on ETMY, I noticed that the open voltages for the UR and LL OSEMs had significantly (>30%) changed from their values from ~2 years ago. The fact that it only occurred in 2 coils seemed to rule out gradual wear and tear, so I looked up the trends from Nov 25 - Nov 28 (Sundance visited on Nov 26 which is when we removed the cage). Not surprisingly, these are the exact two OSEMs that show a decrease in sensor voltage when the OSEMs were pulled out. I suspect that when I placed them in their little Al foil boats, I shorted out some contacts on the rear (this is reminiscent of the problem we had on PRM in 2016). I hope the problem is with the current buffer IC in the satellite box and not the physical diode, I'll test with the tester box and evaluate the problem further.

Attachment 1: Screen_Shot_2019-01-19_at_3.32.35_PM.png  198 kB  | Hide | Hide all
Screen_Shot_2019-01-19_at_3.32.35_PM.png
Attachment 2: ETMY_sensors_1231832635.pdf  31 kB  | Hide | Hide all
ETMY_sensors_1231832635.pdf
ELOG V3.1.3-